Given the following model:

Consumption: C = 500 + 0.5Yd

 Investment: I = 250

Government Expenditure: G = 100

Proportional Tax Rate: t = 0.1

Imports: M = 0.25Y

Exports: X = 50

(Note: There is no lump-sum tax)


  1. a) If the current level of output is 1000, what is the level of actual investment? (2 marks)
  2. b) Calculate the equilibrium real GDP (2 marks)
1
Expert's answer
2020-09-11T09:07:52-0400

Since this is an open economy, the national income "Y" is given as below;


"Y=C+I+G+X-M"

a) Given "Y" as 1000, then actual investment can be obtained as


"Y=1000,\ C=500+0.5Yd,\ G=100, X-M=200,M=0.25X1000=250"

Actual investment

"I=Y-C-G-(X-M)\\
I=1000-500+0.5Yd-100-250"

b) Equilibrium


"Y=300+0.5(Y-0.1) +250+100+250\\
=500+0.5Y-0.05+350+200\\
0.95Y=1050\\
Y_e=1105.26"